The following reaction forms 6.41 L of oxygen at a temperature of 18.7oC and a pressure of 731 torr, what mass (in grams) of KClO3 must have decomposed? Round your answer to three sig figs.

2KClO3 (s) --> 2KCl (s) + 3O2 (g)
1
Expert's answer
2020-04-22T09:16:57-0400

P = 731 torr = 731 torr * (1 atm / 760 torr) = 0.9618 atm

T = 18.7oC = (18.7oC + 273.15) = 291.85 K

R = 0.08206 L atm K-1 mol-1


Solution:

Balanced equation for the given chemical reaction:

2KClO3(s) --> 2KCl(s)+ 3O2(g)

Let's use the ideal gas equation to find the moles of O2 that form.

The ideal gas equation can be expressed as: PV = nRT.

n = PV / RT

Then,

n(O2) = (0.9618 atm * 6.41 L) / (0.08206 L atm K-1 mol-1 * 291.85 K) = 0.2574 moles

Moles of O2 = n(O2) = 0.2574 moles

According to the chemical equation: n(KClO3)/2 = n(O2)/3.

n(KClO3) = 2*n(O2)/3;

n(KClO3) = (2 * 0.2574 moles) / 3 = 0.1716 moles.

Moles of KClO3 = n(KClO3) = 0.1716 moles.


Moles of KClO3 = Mass of KClO3 / Molar mass of KClO3

Molar mass of KClO3 = Mr(KClO3) = 122.55 g/mol.

Finally, we can find the number of grams of KClO3 that reacted:

m(KClO3) = n(KClO3) * Mr(KClO3)

m(KClO3) = (0.1716 moles) * (122.55 g/mol) = 21.0296 g = 21.0 g

m(KClO3) = 21.0 g.


Answer: 21.0 grams of KClO3 must have decomposed.